OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution @sodukogirl13

OpenStudy (anonymous):

Distributive Law for both sides of the equation: -3x - 9 = -3x - 3 - 5 Now continue to solve for x.

OpenStudy (anonymous):

@sodukogirl13 can you help

OpenStudy (anonymous):

C i.e. real no.'s coz x is equation is not dependent on x

OpenStudy (anonymous):

-3x - 9 = -3x - 3 - 5 -3x-9=-3x -8 -3x=-3x-17 -6x= -17 x=6/17

OpenStudy (anonymous):

-3x-9=-3x-3-5 +3x +3x -9=-3-5 -9=-8

OpenStudy (anonymous):

no solution?

OpenStudy (anonymous):

correct

OpenStudy (austinl):

\(-3(x + 3) = -3(x + 1) - 5\) \(-3x-9=-3x-3-5\) \(-3x-9=-3x-8\) \(-3x+3x=-8+9\) \(0\ne1\) No solutions.